# Rethinking Range-View LiDAR Segmentation in Adverse Weather

### v1

@@ Abstract
LiDAR segmentation has emerged as an important task to enrich multimedia experiences and analysis. Range-view-based methods have gained popularity due to their high computational efficiency and compatibility with real-time deployment. However, their generalized performance under adverse weather conditions remains underexplored, limiting their reliability in real-world environments. In this work, we identify and analyze the unique challenges that affect the generalization of range-view LiDAR segmentation in severe weather. To address these challenges, we propose a modular and lightweight framework that enhances robustness without altering the core architecture of existing models. Our method reformulates the initial stem block of standard range-view networks into two branches to process geometric attributes and reflectance intensity separately. Specifically, a Geometric Abnormality Suppression (GAS) module reduces the influence of weather-induced spatial noise, and a Reflectance Distortion Calibration (RDC) module corrects reflectance distortions through memory-guided adaptive instance normalization. The processed features are then fused and passed to the original segmentation pipeline. Extensive experiments on different benchmarks and baseline models demonstrate that our approach significantly improves generalization to adverse weather with minimal inference overhead, offering a practical and effective solution for real-world LiDAR segmentation.
@@ 1 Introduction
LiDAR segmentation has become an essential component in multimedia content understanding [1, 2, 3, 4], as it provides complementary 3D geometric information that traditional 2D visual data typically lacks. By assigning semantic labels (e.g., vehicles, pedestrians, road surfaces) to individual LiDAR points, semantic segmentation helps to enhance the analysis of scene geometry and contents. Recent advancements in deep learning, coupled with the availability of specialized point-cloud semantic segmentation datasets, have driven notable progress in this area [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15]. Despite these advancements, a critical limitation remains: most existing models are tailored for clear weather conditions, and often experience significant performance drops when exposed to adverse environments such as rain, snow, or fog [16, 17, 18, 19]. This vulnerability highlights the urgent need for more robust semantic segmentation methods that can maintain accuracy and reliability across diverse and unpredictable weather conditions, which is a critical requirement for real-world deployment.
@@
To improve generalization under adverse weather, prior efforts have explored physics-based simulation [20, 19, 16, 21] and task-agnostic data augmentation [18, 22, 23]. While these strategies offer promising results, physics-based simulations often fall short in capturing the diverse scope and severity of real-world weather conditions, and augmentation-based approaches may incur extra training overhead and irrelevant variations. Moreover, these methods primarily focus on voxel-based segmentation model [18, 22, 20], overlooking the more efficient range-view-based alternatives. In comparison, range-view methods significantly reduce computational costs and simplify deployment by leveraging well-optimized and widely supported standard operations like 2D convolution, making them especially suitable for latency-sensitive and resource-constrained applications.
@@ 3.1 Preliminaries
We consider the task of domain-generalized LiDAR semantic segmentation. The objective is to train a segmentation model using only source domain data collected under clear-weather conditions, such that it generalizes effectively to unseen target domains characterized by adverse weather (e.g., rain, snow, fog), while maintaining high accuracy on the source domain and incurring minimal additional inference latency. Formally, we are given a labeled source domain dataset $\mathcal{S}=\{(P^{s}_{i},L^{s}_{i})\}_{i=1}^{N^{s}}$, where $N^{s}$ denotes the number of LiDAR scans. Each sample $(P^{s}_{i},L^{s}_{i})$ consists of a point cloud $P^{s}_{i}$ containing $n$ points, and corresponding point-wise semantic labels $L^{s}_{i}$. The point cloud is defined as $P^{s}_{i}=\{(x^{s}_{ij},y^{s}_{ij},z^{s}_{ij},r^{s}_{ij})\}^{n}_{j=1}$, where $(x,y,z)$ denotes the 3D coordinates and $r$ represents the reflectance intensity of each point. The goal is to learn a semantic segmentation model $f:P\rightarrow L$ based solely on the source domain data $\mathcal{S}$, such that it performs well on an unlabeled target domain $\mathcal{T}=\{P^{t}_{i}\}_{i=1}^{N^{t}}$, which remains entirely unseen during training.
@@ 3.2 Geometric Abnormality Suppression
And an Abnormality Network $f_{\theta}(\cdot)$ formulated by several $1\times 1$ convolutional blocks followed by a softmax layer is applied to distinguish between them:
@@ 3.3 Reflectance Distortion Calibration
$\hat{F}_{\text{ref}}=\sigma_{source}(F_{\text{ref}})(\frac{F_{\text{ref}}-\mu_{\text{ref}}}{\sigma_{\text{ref}}})+\mu_{source}(F_{\text{ref}})$ (8)
where $\mu_{\text{ref}}$ and $\sigma_{\text{ref}}$ denote the channel-wise statistics of $F_{\text{ref}}$. Meanwhile, $\mu_{\text{source}}(F_{\text{ref}})$ and $\sigma_{\text{source}}(F_{\text{ref}})$ represent the input-correlated style statistics of the source domain, which are adaptively estimated to achieve robust alignment under varying conditions. To this end, we introduce a parametric Source Style Memory Bank $\mathcal{M}\in\mathbb{R}^{T\times C}$, where $T$ is the number of style vectors and $C$ is the feature dimension. This memory bank stores style vectors learned on the source domain.
@@
$\mu_{\mathrm{source},c}(F_{\text{ref}})$ $=\;\frac{1}{HW}\,\sum_{i=1}^{H}\sum_{w=1}^{W}V_{chw},$ (10)
$\sigma_{\mathrm{source},c}(F_{\text{ref}})$ $=\;\sqrt{\frac{1}{HW}\,\sum_{h=1}^{H}\sum_{w=1}^{W}\Bigl{(}V_{chw}-\mu_{\mathrm{source},c}(F_{\text{ref}})\Bigr{)}^{2}+\epsilon}.$
@@
Style Calibration. After retrieving the source-domain statistics $\mu_{\mathrm{source}}(F_{\text{ref}}^{\text{aug}})$ and $\sigma_{\mathrm{source}}(F_{\text{ref}}^{\text{aug}})$ from the augmented style retrieval step, we perform a final calibration of the original reflectance features. Specifically, we replace the channel-wise mean and variance of $F_{\text{ref}}$ with the learned source statistics, transforming $F_{\text{ref}}$ into a clean-weather style:
$\hat{F}^{\text{aug}}_{\text{ref}}\;=\;\sigma_{\mathrm{source}}(F^{\text{aug}}_{\text{ref}})\left(\frac{F_{\text{ref}}-\mu_{\text{ref}}}{\sigma_{\text{ref}}}\right)\;+\;\mu_{\mathrm{source}}(F^{\text{aug}}_{\text{ref}}),$ (12)
@@
$\begin{split}\mathcal{L}_{SA}\;=\;\bigl{\|}\mu_{\mathrm{source}}\bigl{(}F^{\text{aug}}_{\text{ref}}\bigr{)}-\mu\bigl{(}F_{\text{ref}}\bigr{)}\bigr{\|}~{}~{}~{}\\ \;+\;\bigl{\|}\sigma_{\mathrm{source}}\bigl{(}F^{\text{aug}}_{\text{ref}}\bigr{)}-\sigma\bigl{(}F_{\text{ref}}\bigr{)}\bigr{\|}.\end{split}$ (14)
@@ 4.2 Main Results
SemanticKITTI to SemanticSTF.  As shown in Table 1, our method significantly enhances generalization performance on the SemanticSTF dataset. Compared to baseline range-view models, we observe substantial improvements in mIoU on the testing domain. These gains are largely due to our approach’s ability to suppress spatial noise and reflectance intensity distortion. For example, compared to the original SalsaNext model, our approach shows impressive improvements under different weather conditions: +15.9 mIoU in dense fog, +19.8 mIoU in light fog, +14.5 mIoU in rain, and +18.0 mIoU in snow, with an overall average gain of +20.3 mIoU across all classes. Notably, applying our method to RangeViT results in the best performance among all range-view methods, achieving an increase from 10.2 mIoU to 28.9 mIoU, surpassing even the voxel-based method MinkNet+PDR [18], which achieves 28.6 mIoU. This demonstrates that our method not only excels in improving performance on a variety of weather conditions but also competes favorably against voxel-based approaches. When evaluating the performance on source domain (SemanticKITTI), our method outperforms several baseline range-view models, as shown in the ”Source” column of Table 1. This indicates that our approach effectively maintain accuracy on this source domain while improving performance on the target domain under adverse-weather conditions.
@@
Our method is particularly effective at segmenting classes that are challenging for other methods. For instance, in the ”car” category, the proposed method achieves a substantial improvement of +46.0 mIoU over the baseline SalsaNext. Similarly, we see notable gains in the ”person” category (+41.2 mIoU) and ”road” category (+18.9 mIoU), which are typically difficult to segment accurately under adverse conditions. These improvements highlight our method’s robustness and its ability to handle both common and safety-critical classes that often suffer from performance degradation in existing models. Furthermore, we notice that applying the augmentation techniques in the existing method PDR [18] leads to a performance decrease on CNN-based models, and very minor improvement for the transformer-based model RangeVIT. In contrast, our approach achieves substantial improvements across most semantic categories. Similar performance improvements are observed across other models, further showcasing the broad applicability and effectiveness of our approach. Qualitative results are presented in the supplementary materials.

### v3

# Towards Generalized Range-View LiDAR Segmentation in Adverse Weather

@@ Abstract
LiDAR segmentation has emerged as an important task to enrich scene perception and understanding. Range-view-based methods have gained popularity due to their high computational efficiency and compatibility with real-time deployment. However, their generalized performance under adverse weather conditions remains underexplored, limiting their reliability in real-world environments. In this work, we identify and analyze the unique challenges that affect the generalization of range-view LiDAR segmentation in severe weather. To address these challenges, we propose a modular and lightweight framework that enhances robustness without altering the core architecture of existing models. Our method reformulates the initial stem block of standard range-view networks into two branches to process geometric attributes and reflectance intensity separately. Specifically, a Geometric Abnormality Suppression (GAS) module reduces the influence of weather-induced spatial noise, and a Reflectance Distortion Calibration (RDC) module corrects reflectance distortions through memory-guided adaptive instance normalization. The processed features are then fused and passed to the original segmentation pipeline. Extensive experiments on different benchmarks and baseline models demonstrate that our approach significantly improves generalization to adverse weather with minimal inference overhead, offering a practical and effective solution for real-world LiDAR segmentation.
@@ 1 Introduction
LiDAR segmentation has become an essential component in scene understanding, as it provides complementary 3D geometric information that traditional 2D visual data typically lacks. By assigning semantic labels (e.g., vehicles, pedestrians, road surfaces) to individual LiDAR points, semantic segmentation helps to enhance the analysis of scene geometry and contents. Recent advancements in deep learning, coupled with the availability of specialized point-cloud semantic segmentation datasets, have driven notable progress in this area [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12]. Despite these advancements, a critical limitation remains: most existing models are tailored for clear weather conditions, and often experience significant performance drops when exposed to adverse environments such as rain, snow, or fog [13, 14, 15, 16]. This vulnerability highlights the urgent need for more robust semantic segmentation methods that can maintain accuracy and reliability across diverse and unpredictable weather conditions, which is a critical requirement for real-world deployment.
@@
To improve generalization under adverse weather, prior efforts have explored physics-based simulation [17, 16, 13, 18] and task-agnostic data augmentation [15, 19, 20]. While these strategies offer promising results, physics-based simulations often fall short in capturing the diverse scope and severity of real-world weather conditions, and augmentation-based approaches may incur extra training overhead and irrelevant variations. Moreover, these methods primarily focus on voxel-based segmentation model [15, 19, 17, 21], overlooking the more efficient range-view-based alternatives. In comparison, range-view methods significantly reduce computational costs and simplify deployment by leveraging well-optimized and widely supported standard operations like 2D convolution, making them especially suitable for latency-sensitive and resource-constrained applications.
@@ 3.1 Preliminaries
[…]
@@ 3.2 Geometric Abnormality Suppression
And an Abnormality Network $f_{\theta}(\cdot)$ formulated by several $\times 1$ 1 convolutional blocks followed by a softmax layer is applied to distinguish between them:
@@ 3.3 Reflectance Distortion Calibration
$\hat{F}_{\text{ref}}=\sigma_{src}(F_{\text{ref}})(\frac{F_{\text{ref}}-\mu_{\text{ref}}}{\sigma_{\text{ref}}})+\mu_{src}(F_{\text{ref}})$ (8)
where $\mu_{\text{ref}}$ and $\sigma_{\text{ref}}$ denote the channel-wise statistics of $F_{\text{ref}}$. Meanwhile, $\mu_{\mathrm{src}}(F_{\text{ref}})$ and $\sigma_{\mathrm{src}}(F_{\text{ref}})$ represent the input-correlated style statistics of the source domain, which are adaptively estimated to achieve robust alignment under varying conditions. To this end, we introduce a parametric Source Style Memory Bank $\mathcal{M}\in\mathbb{R}^{T\times C}$, where $T$ is the number of style vectors and $C$ is the feature dimension. This memory bank stores style vectors learned on the source domain.
@@
$\mu_{\mathrm{src},c}(F_{\text{ref}})$ $=\frac{1}{HW}\,\sum_{i=1}^{H}\sum_{w=1}^{W}V_{chw},$ (10)
$\sigma_{\mathrm{src},c}(F_{\text{ref}})$ $=\sqrt{\frac{1}{HW}\,\sum_{h=1}^{H}\sum_{w=1}^{W}(V_{chw}-\mu_{\mathrm{src},c}(F_{\text{ref}}))^{2}}$
@@
Style Calibration. After retrieving the source-domain statistics $\mu_{\mathrm{src}}(F_{\text{ref}}^{\text{aug}})$ and $\sigma_{\mathrm{src}}(F_{\text{ref}}^{\text{aug}})$ from the augmented style retrieval step, we perform a final calibration of the original reflectance features. Specifically, we replace the channel-wise mean and variance of $F_{\text{ref}}$ with the learned source statistics, transforming $F_{\text{ref}}$ into a clean-weather style:
$\hat{F}^{\text{aug}}_{\text{ref}}\;=\;\sigma_{\mathrm{src}}(F^{\text{aug}}_{\text{ref}})\left(\frac{F_{\text{ref}}-\mu_{\text{ref}}}{\sigma_{\text{ref}}}\right)\;+\;\mu_{\mathrm{src}}(F^{\text{aug}}_{\text{ref}}),$ (12)
@@
$\begin{split}\mathcal{L}_{SA}\;=\;\bigl{\|}\mu_{\mathrm{src}}\bigl{(}F^{\text{aug}}_{\text{ref}}\bigr{)}-\mu\bigl{(}F_{\text{ref}}\bigr{)}\bigr{\|}~~~\\ \;+\;\bigl{\|}\sigma_{\mathrm{src}}\bigl{(}F^{\text{aug}}_{\text{ref}}\bigr{)}-\sigma\bigl{(}F_{\text{ref}}\bigr{)}\bigr{\|}.\end{split}$ (14)
@@ 4.2 Main Results
SemanticKITTI to SemanticSTF.  As shown in Table 1, our method significantly enhances generalization performance on the SemanticSTF dataset. Compared to baseline range-view models, we observe substantial improvements in mIoU on the testing domain. These gains are largely due to our approach’s ability to suppress spatial noise and reflectance intensity distortion. For example, compared to the original SalsaNext model, our approach shows impressive improvements under different weather conditions: +15.9 mIoU in dense fog, +19.8 mIoU in light fog, +14.5 mIoU in rain, and +18.0 mIoU in snow, with an overall average gain of +20.3 mIoU across all classes. Notably, applying our method to RangeViT results in the best performance among all range-view methods, achieving an increase from 10.2 mIoU to 28.9 mIoU, surpassing even the voxel-based method MinkNet+PDR [15], which achieves 28.6 mIoU. This demonstrates that our method not only excels in improving performance on a variety of weather conditions but also competes favorably against voxel-based approaches. When evaluating the performance on source domain (SemanticKITTI), our method outperforms several baseline range-view models, as shown in the ”Source” column of Table 1. This indicates that our approach effectively maintain accuracy on this source domain while improving performance on the target domain under adverse-weather conditions. We further evaluated the current state-of-the-art voxel-based method GRC [47], adapted to the range-view setting, for comparison with our approach. Experimental results show that our method still shows significant advantages on SalsaNext compared to GRC, with mIoU increasing by +6.3. The same advantages are also shown on other models. It is worth noting that both RDA [19] and GRC [47] exhibit varying degrees of performance degradation on the source domain, further highlighting the advantages of our approach.
@@
Our method is particularly effective at segmenting classes that are challenging for other methods. For instance, in the ”car” category, the proposed method achieves a substantial improvement of +46.0 mIoU over the baseline SalsaNext. Similarly, we see notable gains in the ”person” category (+41.2 mIoU) and ”road” category (+18.9 mIoU), which are typically difficult to segment accurately under adverse conditions. These improvements highlight our method’s robustness and its ability to handle both common and safety-critical classes that often suffer from performance degradation in existing models. Furthermore, we notice that applying the augmentation techniques in the existing method PDR [15] leads to a performance decrease on CNN-based models, and very minor improvement for the transformer-based model RangeVIT. In contrast, our approach achieves substantial improvements across most semantic categories. Similar performance improvements are observed across other models, further showcasing the broad applicability and effectiveness of our approach. Some qualitative results under different weathers are presented in the Figure  4.
In addition, we further verified the effectiveness of our method on voxel-based backbone. As shown in Figure  3, we conducted a preliminary study by applying our RDC and GAS modules to a voxel-based MinkNet, and achieved 32.4 mIoU. Compared to the simple MinkNet, this represents an improvement of +7.8 mIoU. Our approach also outperform the baseline in various weather types. While these results do not yet surpass current voxel-based SOTA methods, they demonstrate substantial performance improvements, suggesting that the core ideas behind RDC and GAS are indeed generalizable beyond the range-view setting.
@@
Figure 3: Generalized segmentation performance with MinkNet as backbone on SemanticSTF.
Figure 4: Qualitative results of our method, trained on the SemanticKITTI training set, evaluated under the light fog, dense fog, rain and snow conditions on the SemanticSTF validation set. Green points denote correct predictions, whereas red points indicate incorrect predictions.